Q.The remainder when 11997 +21997 +31997+..........................19961997  is divided by 1997 is-


a.0       b.1   c.197     d.1996


(Plz provide the procedure as well)

Grade:9

3 Answers

Akash Kumar Dutta
10 years ago

by applying the formula a^n + b^n= (a+b)(a^n-1 - a^n-2.b.....+b^n)
we group the terms 1,1996  .... 2, 1995 .....so on
by this the common factor comes as 1+1996 or 2+1995 ...=1997
hence 1997 perfectly divides the above value and remainder = 0
Plz Approve

Akash Kumar Dutta
10 years ago

i have already given the solution of this question since it has been repeated .
please follow that and do approve,
by the way the asn. is 0

Abhishekh kumar sharma
10 years ago

0